OpenStudy (anonymous):

A state vector X for a four-state Markov chain is such that the system is four times as likely to be in state 3 as in 4, is not in state 2, and is in state 1 with probability 0.2. Find the state vector X.

OpenStudy (unklerhaukus):

P(1) + P(2) + P(3) + P(4) = 1 0.2 + 0 + 4P(4) + P(4) = 1

OpenStudy (unklerhaukus):

solve for P(4)

OpenStudy (unklerhaukus):

\[\mathbf X = \begin{bmatrix}P(1)\\P(2)\\P(3)\\P(4)\end{bmatrix}\]